OpenStudy (anonymous):

Hello, I am new to power series calculus and need help in proving the following implication (in the next post).

OpenStudy (anonymous):

If \[\sum_{n=0}^{\infty}c_n 4^n\] is convergent, does it follow that the following series are convergent? a) \[\sum_{n=0}^{\infty}c_n(-2)^n\] b) \[\sum_{n=0}^{\infty}c_n(-4)^n\]

OpenStudy (anonymous):

lol